OpenStudy (samsan9):

given f(x)=1/x+3. Find the domain and range

OpenStudy (samsan9):

\[f(x)= \frac{ 1 }{ x+3 }\]

OpenStudy (dumbcow):

Domain: find where the function in undefined .... set denominator equal to 0 x+3 = 0 x = -3 so domain is all real numbers except -3 Range: the denominator has higher degree so the horizontal asymptote is y=0 Range is all real numbers except 0

OpenStudy (samsan9):

thank you and do you think you could me with another ? i already have the domain i just need the range?

OpenStudy (dumbcow):

ok

OpenStudy (samsan9):

\[\sqrt{x+6}\]

OpenStudy (dumbcow):

the sqrt of a number is always positive Range is y>= 0

OpenStudy (samsan9):

so how do you write it?

OpenStudy (dumbcow):

lol y>= 0 or in interval form ..... [0, infinity)

OpenStudy (samsan9):

lol yeah i forgot what it was called but thanks XD

OpenStudy (dumbcow):

yw